OpenStudy (anonymous):

K is the midpoint of . The coordinates of K are (3, 5) and the coordinates of L are (-5, -6), find the coordinates of M.

OpenStudy (anonymous):

Is K the Midpoint of LM?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

 \[3=\frac{ -5+x }{2 },-5+x=6,x=6+5=11\] \[5=\frac{ -6+y }{2 },-6+y=10,y=10+6=16\]

OpenStudy (anonymous):

so what would the answer be exactly

OpenStudy (anonymous):

M(11,16) You can check this using the midpoint formula